OpenStudy (anonymous):

Solve the equation: 4x2 = 9

OpenStudy (anonymous):

4x2= 8

OpenStudy (anonymous):

i'm thinking your equation is this \[4x ^{2}=9\]. so x^2 = 9/4 x = 3/2 or -3/2